### v2

@@ 3.1 The reduction
We now are ready to argue that $G$ is $P_{14}$-free. Suppose not and let $P:v_{1},v_{2},\allowbreak\ldots,v_{14}$ be an induced $P_{14}$ in $G$, with edges $v_{i}v_{i+1}$, $141\leq i<$ 14. For $i<j$, write $P[v_{i},v_{j}]$ for the subpath of $P$ between (and including) $v_{i}$ and $v_{j}$. Then, by (2), each of $P[v_{1},v_{7}]$ and $P[v_{8},v_{14}]$ contains a vertex from the clique $T$. Since $P$ has no chords, $P[v_{1},v_{7}]$ has only the vertex $v_{7}$ in $T$ and $P[v_{8},v_{14}]$ has only the vertex $v_{8}$ in $T$. By (1), therefore, both $P[v_{1},v_{6}]$ and $P[v_{9},v_{14}]$ contain some vertex in the clique $F$, and thus $P$ has a chord. This contradiction shows that $G$ is $P_{14}$-free, as claimed.
We now show that $G$ is $8$-chordal. First observe that, if $S$ is a clique cutset in (an arbitrary graph) $G$ then any induced cycle in $G$ is contained in $G[A\cup S]$ for some connected component $A$ of $G-S$. With this observation, we have
$G-T$ has no induced cycles $C_{k}$ of length $k\geq 7$. (3)
This can be seen as follows: write $F=C\cup C^{\prime}$ with $C=\{c_{j}\mid 1\leq j\leq m\}$ and $C^{\prime}=\{c_{j}^{\prime}\mid 1\leq j\leq m\}$. Then $C$ is a clique cutset of $G-T$ separating each variable clique $Q(x)$ and the connected component $H$ consisting of $C^{\prime}$ and all clause $6$-cycles $D_{j}$. Hence each induced cycle in $G-T$ is contained in some $Q(x)\cup C$ or in $H\cup C$. Since $Q(x)\cup C$ induces a co-bipartite graph, it has no induced cycle of length larger than $4$. Furthermore, $C^{\prime}$ is a clique cutset of the graph induced by $H\cup C$ separating each $D_{j}$ and $C$. Hence each induced cycle in $H\cup C$ is contained in some $D_{j}\cup C^{\prime}$ or in $C\cup C^{\prime}=F$. In each $D_{j}\cup C^{\prime}$, $c_{j}^{\prime}$ is the cut-vertex separating $D_{j}$ and $C^{\prime}\setminus\{c_{j}^{\prime}\}$. Since $c_{j}^{\prime}$ is adjacent to three vertices of the $6$-cycle $D_{j}$ and $F$ is a clique, $D_{j}\cup C^{\prime}$ and hence $H\cup C^{\prime}$ has no induced cycles of length larger than $6$. Thus, (3) is proved.
Now, suppose for the contrary that $G$ has an induced cycle $Z$ of length at least  $9$. Then, by (3), $Z\cap T\not=\emptyset$. Since $T$ is a clique, $Z\cap T$ is a vertex or two adjacent vertices. But then $Z-T$ is an induced path in $G-T$ with at leat $7$ vertices, contradicting (2). Thus, $G$ is $8$-chordal. $\qed$
@@
    if $F\subseteq X$, then |$|\{c_{j1},c_{j2},c_{j3}\}\cap Y|=1$, and if $F\subseteq Y$, then |$|\{c_{j1},c_{j2},\allowbreak c_{j3}\}\cap X|=1$;
@@
    if $F\subseteq X$, then |$|\{b_{j1},b_{j2},b_{j3}\}\cap Y|=2$ and |$|\{c_{j1}^{\prime},c_{j2}^{\prime},c_{j3}^{\prime}\}\cap Y|=1$, and if $F\subseteq Y$, then |$|\{b_{j1},b_{j2},b_{j3}\}\cap X|=2$ and |$|\{c_{j1}^{\prime},c_{j2}^{\prime},c_{j3}^{\prime}\}\cap X|=1$.
@@
Similarly, it can be seen that every vertex in $Y$ has a neighbor in $X$$\qed$
@@
If $\phi$ has an 1-in-3 assignment, then $G$ has a perfect matching cut.
@@
[…]
@@
Let $Y=V(G)\setminus X$. Then, it is not difficult to verify that $M=(X,Y)$ is a perfect matching cut of $G$$\qed$
$T$$F$$F$$c_{1}$$x$$y$$z$$a_{11}$$a_{12}$$a_{13}$$b_{11}$$b_{12}$$b_{13}$$c_{11}^{\prime}$$c_{12}^{\prime}$$c_{13}^{\prime}$$c_{1}^{\prime}$$c_{2}$$u$$z$$y$$a_{21}$$a_{22}$$a_{23}$$b_{21}$$b_{22}$$b_{23}$$c_{21}^{\prime}$$c_{22}^{\prime}$$c_{23}^{\prime}$$c_{2}^{\prime}$$c_{3}$$z$$v$$w$$a_{31}$$a_{32}$$a_{33}$$b_{31}$$b_{32}$$b_{33}$$c_{31}^{\prime}$$c_{32}^{\prime}$$c_{33}^{\prime}$$c_{3}^{\prime}$
@@
Now, suppose $\phi$ has an 1-in-3 assignment. Then, by Lemma 4, $G$ has a perfect matching cut. In particular, $G$ has a disconnected perfect matching and, actually, a matching cut.
Conversely, let $G$ have a matching cut $M=(X,Y)$, possibly a perfect matching cut or one that is contained in a perfect matching of $G$. Then, by Lemma 2 (i), we may assume that $F\subseteq X$, and set variable $x$ to True if the corresponding variable clique $Q(x)$ is contained in $Y$ and False if $Q(x)$ is contained in $X$. By Lemma 2 (iii), this assignment is well defined. Moreover, it is an 1-in-3 assignment for $\phi$: consider a clause $C_{j}=\{x,y,z\}$ with $c_{j1}=x,c_{j2}=y$ and $c_{j3}=z$. By Lemma 2 (ii) and (iii), exactly one of $Q(x),Q(y)$ and $Q(z)$ is contained in $Y$, hence exactly one of $x,y$ and $z$ is assigned True.
@@
The proof of Theorem 5 is complete.$\qed$
4 Proof of Theorem 4
Recall that by Theorem 2, mc is polynomially solvable for $4$-chordal graphs (also called quadrangulated graphs). In this section, we first point out that dpm is polynomially solvable for $4$-chordal graphs, too, by following known approaches KratschL16 ; LeL19 ; Moshi89 . We note that the approach in KratschL16 ; LeL19 has some similarity to Moshi’s algorithm Moshi89 for mc on $4$-chordal graphs. However, it is not immediately how to modify the latter for solving dpm restricted to $4$-chordal graphs. We note also that both approaches for mc and dpm do not work for pmc. Then, we will show that pmc can be solved in polynomial time on $4$-chordal graphs by a new approach involving a non-trivial reduction to 2sat.
@@ 4 Proof of Theorem 4
For each edge $ab$ of a $4$-chordal graph $G$, we will be able to decide if $G$ has a disconnected perfect matching containing a matching cut separating $A=\{a\}$ and $B=\{b\}$. This is done by applying known forcing rules (KratschL16 ; LeL19 ), which are given below. Initially, set $X:=A$, $Y:=B$ and write $F=V(G)\setminus(X\cup Y)$ for the set of ‘free’ vertices. The sets $A,B,X$ and $Y$ will be extended, if possible, by adding vertices from $F$ so that $A\subseteq X$, $B\subseteq Y$ and if $(U,W)$ is any $A$,$B$-matching cut of $G$ then either $X\subseteq U$ and $Y\subseteq Y$, or else $X\subseteq W$ and $Y\subseteq X$. In doing so, we will apply the following forcing rules exhaustively. The first three rules will detect certain vertices that ensure that $G$ cannot have an $A,B$-matching cut.
@@
We refer to LeL19 for the correctness of rules (R4) and (R5), and for the following facts.
@@
We now are ready to prove Theorem 4. Let $G$ be a connected, $4$-chordal graph, and let $ab$ be an edge of $G$. Set $A=\{a\}$ and $B=\{b\}$, and assume that none of (R1) – (R5) is applicable. Then, denoting $N(S)$ the set of vertices outside $S$ adjacent to some vertex in $S$, for any connected component $S$ of $G[F]$,
@@
For, otherwise choose two vertices $s,s^{\prime}\in S$ with a neighbor $x\in N(s)\cap X$ and a neighbor $y\in N(s^{\prime})\cap Y$ such that the distance between $s$ and $s^{\prime}$ in $S$ is as small as possible. Then $s,s^{\prime},x$ and $y$ and a shortest $s,s^{\prime}$-path in $S$, a chordless $x,y$-path in $G[X\cup Y]$ together would induce a long hole in $G$. (Observe that, by the definition of $X$ and $Y$, $G[X\cup Y]$ is connected.)
@@
cut if and only if $G-A-B$ has a perfect matching.
Therefore, with Fact 1, in $O(nm)$ time we can decide whether $G$ has a matching cut containing a given edge. [^3] Thus, mc is solvable in $O(nm^{2})$ time when restricted to $4$-chordal graphs, an alternative proof of Theorem 2. We remark that Moshi’s algorithm solving mc in $4$-chordal graphs runs in $O(n^{3}m)$ time which is worse than ours for sparse graphs. Moreover, as a maximum matching can be computed in $O(\sqrt{n}m)$ time MicaliV80 , we can decide in $O(n\sqrt{n}m^{2})$ time whether $G$ has a disconnected perfect matching containing an $\{a\},\{b\}$-matching cut for a given edge $ab$. Since there are at most $m$ edges to check, we can solve dpm in $O(n\sqrt{n}m^{3})$ time on $4$-chordal graphs.
We note that the above approach cannot be adapted to solve pmc because the last fact is not longer true for perfect matching cuts. For instance, if $G-A-B$ has a perfect matching cut then $G$ does not need to have a perfect matching cut containing an $A,B$-matching cut. In the next subsection, we will reduce pmc restricted to $4$-chordal graphs to 2sat instead.
@@
Let $h$ be the largest integer such that $L_{h}\not=\emptyset$ but $L_{h+1}=\emptyset$. Note that $L_{0}=\{r\}$ and that all level sets $L_{i}$, $0\leq i\leq h$, can be computed in linear time by running a breadth-first-search at $r$. To avoid triviality, we assume that $h\geq 2$.
@@
Now, let $S\subseteq N(v)\cap L_{i-1}$ be an independent set and consider a vertex $w\in L_{i-2}$ with maximum $|N(w)\cap S|$. Then $N(w)\cap S=S$, that is, $w$ is adjacent to all vertices in $S$. For otherwise, let $u_{1}\in S\setminus N(w)$ and let $w^{\prime}\in L_{i-2}$ be a neighbor of $u_{1}$. By the choice of $w$, $w^{\prime}$ is non-adjacent to a vertex $u_{2}\in N(w)\cap S$. But then the two vertices $u_{1}$ and $u_{2}$ contradict the statement above.
Recall that a vertex set $S\subseteq V$ is monochromatic if, for every matching cut $(X,Y)$ of $G$$S$ is contained in $X$ or else in $Y$. Observe that if $G[S]$ is connected then $S$ is monochromatic if and only if each edge $\{x,y\}$ of $G[S]$ is monochromatic.
@@
We proceed by induction on $i$. Let $A$ be a connected component in $G[L_{1}]$ with at least two vertices. Since all vertices in $A$ are adjacent to $r$, every edge in $A$ forms with $r$ a triangle. Hence each edge in $A$ is monochromatic, and the statement is true for $i=1$.
Let us assume that the statement is true for $i<h$, and consider a connected component $A$ of $G[L_{i+1}]$. Let $\{x,y\}$ be an edge of $A$ which is not contained in a triangle (otherwise $\{x,y\}$ is monochromatic and we are done), and let $x^{\prime}$, respectively, $y^{\prime}$ be a neighbor in $L_{i}$ of $x$, respectively, $y$. Then $x^{\prime}$ is not adjacent to $y$ and $y^{\prime}$ is not adjacent to $x$; in particular, $x^{\prime}\not=y^{\prime}$. Now, $x^{\prime}$ and $y^{\prime}$ must be adjacent, otherwise a shortest path in $G[L_{0}\cup L_{1}\cup\cdots\cup L_{i-1}\cup\{x^{\prime},y^{\prime}\}]$ connecting $x^{\prime}$ and $y^{\prime}$ together with $x$ and $y$ would form an induced cycle of length at least $5$ in $G$. Now, by induction hypothesis, $\{x^{\prime},y^{\prime}\}$ is monochromatic, implying $\{x,y\}$ is monochromatic, too. $\qed$
@@
Now, to decide whether $G$ admits a perfect matching cut we proceed as follows. First, we create for each vertex $v$ of $G$ a Boolean variable, which also denoted $v$ as the context will be clear. Then, at the beginning, all vertices are undetermined, $Q=\emptyset$, and we process bottom-up with a leaf $v$ in $L_{h}$ by checking if $v$ satisfies one of the conditions in Lemma 7. If not, $G$ has no perfect matching cut. Otherwise, we create certain 2-cnf clauses that assign $v$ to its private neighbor $p(v)$, and force $v$ and $x$, and $p(v)$ and $y$ to be monochromatic for all other neighbors $x\not=p(v)$ of $v$ and all other neighbors $y\not=v$ of $p(v)$. Then $v$ and its private neighbor $p(v)$ become determined. This way we will correctly decide if $G$ has no perfect matching cut or we will successfully construct a 2sat-instance $\phi$ such that $G$ has a perfect matching cut if and only if $\phi$ is satisfiable.
@@
1 fix a vertex $r\in V$ and compute the BFS-levels $L_{i}$ from $r$, $0\leq i\leq h$ create for each vertex $v\in V$ a Boolean variable $v$ $Q\leftarrow\emptyset$; $\phi\leftarrow\emptyset$ for $i\leftarrow h$ downto $1$ do

### v3

@@ 3.1 The reduction
We now are ready to argue that $G$ is $P_{14}$-free. Suppose not and let $P:v_{1},v_{2},\allowbreak\ldots,v_{14}$ be an induced $P_{14}$ in $G$, with edges $v_{i}v_{i+1}$, $1\leq i<14$. For $i<j$, write $P[v_{i},v_{j}]$ for the subpath of $P$ between (and including) $v_{i}$ and $v_{j}$. Then, by (2), each of $P[v_{1},v_{7}]$ and $P[v_{8},v_{14}]$ contains a vertex from the clique $T$. Since $P$ has no chords, $P[v_{1},v_{7}]$ has only the vertex $v_{7}$ in $T$ and $P[v_{8},v_{14}]$ has only the vertex $v_{8}$ in $T$. By (1), therefore, both $P[v_{1},v_{6}]$ and $P[v_{9},v_{14}]$ contain some vertex in the clique $F$, and thus $P$ has a chord. This contradiction shows that $G$ is $P_{14}$-free, as claimed.
We now show that $G$ is $8$-chordal. First observe that, if $S$ is a clique cutset in (an arbitrary graph) $G$ then any induced cycle in $G$ is contained in $G[A\cup S]$ for some connected component $A$ of GSG-S. With this observation, we have
GTG-T has no induced cycles $C_{k}$ of length $k\geq 7$. (3)
This can be seen as follows: write $F=C\cup C^{\prime}$ with $C=\{c_{j}\mid 1\leq j\leq m\}$ and $C^{\prime}=\{c_{j}^{\prime}\mid 1\leq j\leq m\}$. Then $C$ is a clique cutset of GTG-T separating each variable clique $Q(x)$ and the connected component $H$ consisting of $C^{\prime}$ and all clause $6$-cycles $D_{j}$. Hence each induced cycle in GTG-T is contained in some $Q(x)\cup C$ or in $H\cup C$. Since $Q(x)\cup C$ induces a co-bipartite graph, it has no induced cycle of length larger than $4$. Furthermore, $C^{\prime}$ is a clique cutset of the graph induced by $H\cup C$ separating each $D_{j}$ and $C$. Hence each induced cycle in $H\cup C$ is contained in some $D_{j}\cup C^{\prime}$ or in $C\cup C^{\prime}=F$. In each $D_{j}\cup C^{\prime}$, $c_{j}^{\prime}$ is the cut-vertex separating $D_{j}$ and $C^{\prime}\setminus\{c_{j}^{\prime}\}$. Since $c_{j}^{\prime}$ is adjacent to three vertices of the $6$-cycle $D_{j}$ and $F$ is a clique, $D_{j}\cup C^{\prime}$ and hence $H\cup C^{\prime}$ has no induced cycles of length larger than $6$. Thus, (3) is proved.
Now, suppose for the contrary that $G$ has an induced cycle $Z$ of length at least  99. Then, by (3), $Z\cap T\not$$\emptyset$. Since $T$ is a clique, $Z\cap T$ is a vertex or two adjacent vertices. But then ZTZ-T is an induced path in GTG-T with at leat 77 vertices, contradicting (2). Thus, $G$ is $8$-chordal. $\qed$
@@
    if $F\subseteq X$, then $|\{c_{j1},c_{j2},c_{j3}\}\cap Y|=1$, and if $F\subseteq Y$, then $|\{c_{j1},c_{j2},\allowbreak c_{j3}\}\cap X|=1$;
@@
    if $F\subseteq X$, then $|\{b_{j1},b_{j2},b_{j3}\}\cap Y|=2$ and $|\{c_{j1}^{\prime},c_{j2}^{\prime},c_{j3}^{\prime}\}\cap Y|=1$, and if $F\subseteq Y$, then $|\{b_{j1},b_{j2},b_{j3}\}\cap X|=2$ and $|\{c_{j1}^{\prime},c_{j2}^{\prime},c_{j3}^{\prime}\}\cap X|=1$.
@@
[…]
@@
If $\phi$ has a 1-in-3 assignment, then $G$ has a perfect matching cut.
@@
Partition $V(G)$ into disjoint subsets $X$ and $Y$ as follows. (Fig. 5 shows the partition for the example graph in Fig. 4 given the assignment $\allowbreak y=v=\text{True}$, $x=z=u=w=\text{False}$.) First,
@@
[…]
@@
Now, suppose $\phi$ has a 1-in-3 assignment. Then, by Lemma 4, $G$ has a perfect matching cut. In particular, $G$ has a disconnected perfect matching and, actually, a matching cut.
Conversely, let $G$ have a matching cut $M=(X,Y)$, possibly a perfect matching cut or one that is contained in a perfect matching of $G$. Then, by Lemma 2 (i), we may assume that $F\subseteq X$, and set variable $x$ to True if the corresponding variable clique $Q(x)$ is contained in $Y$ and False if $Q(x)$ is contained in $X$. By Lemma 2 (iii), this assignment is well defined. Moreover, it is a 1-in-3 assignment for $\phi$: consider a clause $C_{j}=\{x,y,z\}$ with $c_{j1}=x,c_{j2}=y$ and $c_{j3}=z$. By Lemma 2 (ii) and (iii), exactly one of $Q(x),Q(y)$ and $Q(z)$ is contained in $Y$, hence exactly one of $x,y$ and $z$ is assigned True.
@@
[…]
Remark
It can easily be verified that the constructed graph $G$ has diameter $4$ and radius $3$. There exist a number of results on algorithms and complexity of matching cut problems restricted to graphs of bounded diameter and radius; the paper [23] is an excellent source for this topic. While complexity dichotomies for mc on graphs of bounded diameter and radius ([17, 22]), for pmc on graphs of bounded radius ([13, 22]), and for dpm on graphs of bounded diameter ([4]) are known, it remain only two open cases in this research direction (see [23]): what is the complexity of pmc on graphs of diameter 3, and what is the complexity of dpm on graphs of radius 2?
4 Polynomial results: Proof of Theorem 4
In this section, we prove Theorem 4: dpm and pmc can be solved in polynomial time when restricted to $4$-chordal graphs.
Recall that by Theorem 2, mc is polynomially solvable for $4$-chordal graphs (also called quadrangulated graphs). We first point out that dpm is polynomially solvable for $4$-chordal graphs, too, by following known approaches [16, 17, 26]. We note that the approach in [16, 17] has some similarity to Moshi’s algorithm [26] for mc on $4$-chordal graphs. However, it is not immediately clear how to modify the latter for solving dpm restricted to $4$-chordal graphs. We note also that both approaches for mc and dpm do not work for pmc. Then, we will show that pmc can be solved in polynomial time on $4$-chordal graphs by a new approach involving a non-trivial reduction to 2sat.
@@ 4 Polynomial results: Proof of Theorem 4
For each edge $ab$ of a $4$-chordal graph $G$, we will be able to decide if $G$ has a disconnected perfect matching containing a matching cut separating $A=\{a\}$ and $B=\{b\}$. This is done by applying known forcing rules ([16, 17, 19]), which are given below. Initially, set $X:=A$, $Y:=B$ and write $F=V(G)\setminus(X\cup Y)$ for the set of ‘free’ vertices. The sets $A,B,X$ and $Y$ will be extended, if possible, by adding vertices from $F$ so that $A\subseteq X$, $B\subseteq Y$ and if $(U,W)$ is any AA,$B$-matching cut of $G$ then either $X\subseteq U$ and $Y\subseteq Y$, or else $X\subseteq W$ and $Y\subseteq X$. In doing so, we will apply the following forcing rules exhaustively. The first three rules will detect certain vertices that ensure that $G$ cannot have an $A,B$-matching cut.
@@
We refer to [17] for the correctness of rules (R4) and (R5), and for the following facts.
@@
We now are ready to prove Theorem 4. Let $G$ be a connected, $4$-chordal graph, and let $ab$ be an edge of $G$. Set $A=\{a\}$ and $B=\{b\}$, and assume that none of (R1) – (R5) is applicable. For a connected component $S$ of $G[F]$, let $N(S)$ be the set of vertices outside $S$ adjacent to some vertex in $S$. Then, for any connected component $S$ of $G[F]$,
@@
For, otherwise choose two vertices $s,s^{\prime}\in S$ with a neighbor $x\in N(s)\cap X$ and a neighbor $y\in N(s^{\prime})\cap Y$ such that the distance between $s$ and $s^{\prime}$ in $S$ is as small as possible, as well as the distance between $x$ and $y$ in $G[X\cup Y]$ is as small as possible. Note that, by the choice of $s$ and $s^{\prime}$, if $s$ and $s^{\prime}$ are both adjacent to $x$ or $y$ then $s=s^{\prime}$. Note also that, by the definition of $X$ and $Y$, $G[X\cup Y]$ is connected, and the distance between $x$ and $y$ in $G[X\cup Y]$ is at least three. Then $s,s^{\prime},x$ and $y$ and a shortest $s,s^{\prime}$-path in $S$, a shortest $x,y$-path in $G[X\cup Y]$ together would induce a long hole in $G$.
@@
cut if and only if GABG-A-B has a perfect matching.
Therefore, with Fact 1, in $O(nm)$ time we can decide whether $G$ has a matching cut containing a given edge. [^3] Thus, mc is solvable in $O(nm^{2})$ time when restricted to $4$-chordal graphs, an alternative proof of Theorem 2. We remark that Moshi’s algorithm solving mc in $4$-chordal graphs runs in $O(n^{3}m)$ time which is worse than ours for sparse graphs. Moreover, as a maximum matching can be computed in $O(\sqrt{n}m)$ time [24, 28], we can decide in $O(n\sqrt{n}m^{2})$ time whether $G$ has a disconnected perfect matching containing an $\{a\},\{b\}$-matching cut for a given edge $ab$. Since there are at most $m$ edges to check, we can solve dpm in $O(n\sqrt{n}m^{3})$ time on $4$-chordal graphs.
We note that the above approach cannot be adapted to solve pmc because the last fact is not longer true for perfect matching cuts. For instance, if GABG-A-B has a perfect matching cut then $G$ does not need to have a perfect matching cut containing an $A,B$-matching cut. In the next subsection, we will reduce pmc restricted to $4$-chordal graphs to 2sat instead.
@@
Let $h$ be the largest integer such that $L_{h}\not$$\emptyset$ but $L_{h+1}=\emptyset$. Note that $L_{0}=\{r\}$ and that all level sets $L_{i}$, $0\leq i\leq h$, can be computed in linear time by running a breadth-first-search at $r$. To avoid triviality, we assume that $h\geq 2$. (If $h=1$ then $r$ is a universal vertex. Graphs with at least three vertices and a universal vertex have no perfect matching cut.)
@@
Now, let $S\subseteq N(v)\cap L_{i-1}$ be an independent set and consider a vertex $w\in L_{i-2}$ with maximum $|N(w)\cap S|$. Then $N(w)\cap S=S$, that is, $w$ is adjacent to all vertices in $S$. For otherwise, let $u_{1}\in S\setminus N(w)$ and let $w^{\prime}\in L_{i-2}$ be a neighbor of $u_{1}$. By the choice of $w$, $w^{\prime}$ is non-adjacent to a vertex $u_{2}\in N(w)\cap S$. But then the two vertices $u_{1}$ and $u_{2}$ contradict the statement above. $\qed$
Recall that a vertex set $S\subseteq V$ is monochromatic if, for every matching cut $(X,Y)$ of GGSS is contained in $X$ or else in $Y$. Observe that if $G[S]$ is connected then $S$ is monochromatic if and only if each edge $\{x,y\}$ of $G[S]$ is monochromatic.
@@
We proceed by induction on $i$. Let $A$ be a connected component in $G[L_{1}]$ with at least two vertices. Since all vertices in $A$ are adjacent to the root $r$, every edge in $A$ forms with $r$ a triangle. Hence each edge in $A$ is monochromatic, and the statement is true for $i=1$.
Let us assume that the statement is true for $i<h$, and consider a connected component $A$ of $G[L_{i+1}]$. Let $\{x,y\}$ be an edge of $A$ which is not contained in a triangle (otherwise $\{x,y\}$ is monochromatic and we are done), and let $x^{\prime}$, respectively, $y^{\prime}$ be a neighbor in $L_{i}$ of $x$, respectively, $y$. Then $x^{\prime}$ is not adjacent to $y$ and $y^{\prime}$ is not adjacent to $x$; in particular, $x^{\prime}\not=y^{\prime}$. Now, $x^{\prime}$ and $y^{\prime}$ must be adjacent, otherwise a shortest path in $G[L_{0}\cup L_{1}\cup\cdots\cup L_{i-1}\cup\{x^{\prime},y^{\prime}\}]$ connecting $x^{\prime}$ and $y^{\prime}$ together with $x$ and $y$ would form an induced cycle of length at least $5$ in $G$. Now, by induction hypothesis, $\{x^{\prime},y^{\prime}\}$ is monochromatic, implying $\{x,y\}$ is monochromatic, too. (Otherwise, $G$ has no (perfect) matching cut; $x$ or $y$ would have two neighbors in the other part.) $\qed$
@@
Now, to decide whether $G$ admits a perfect matching cut we proceed as follows. First, we create for each vertex $v$ of $G$ a Boolean variable, which is also denoted by $v$ as the context will be clear. Then, at the beginning, all vertices are undetermined, $Q=\emptyset$, and we process bottom-up with a leaf $v$ in $L_{h}$ by checking if $v$ satisfies one of the conditions in Lemma 7. If not, $G$ has no perfect matching cut. Otherwise, we create certain 2-cnf clauses that assign $v$ to its private neighbor $p(v)$, and force $v$ and $x$, and $p(v)$ and $y$ to be monochromatic for all other neighbors $x\not=p(v)$ of $v$ and all other neighbors $y\not=v$ of $p(v)$. Then $v$ and its private neighbor $p(v)$ become determined. This way we will correctly decide if $G$ has no perfect matching cut or we will successfully construct a 2sat-instance $\phi$ such that $G$ has a perfect matching cut if and only if $\phi$ is satisfiable.
@@
1
21ex
3
4fix a vertex $r\in V$ and compute the BFS-levels $L_{i}$ from $r$, $0\leq i\leq h$
5 create for each vertex $v\in V$ a Boolean variable $v$
6 $Q\leftarrow\emptyset$; $\phi\leftarrow\emptyset$
7
8for $i\leftarrow h$ downto 11 do